OpenStudy (anonymous):

Find the differential y = ln*sqrt(1+x^2)

OpenStudy (anonymous):

\[y=\ln \sqrt{1+x ^{2}}\]

OpenStudy (anonymous):

can u help?

OpenStudy (mertsj):

x/(x^2+1)

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

think it should be \[\frac{2x}{x^2+1}\] by the chain rule

OpenStudy (anonymous):

i know the chain rule is der of first*sec + 1st* der of 2nd but how did u get taht as ur derivative

OpenStudy (mertsj):

derivative of lnx = 1/x so derivative of lnsqrt(1+x^2) is 1/sqrt(1+x^2). But then you have to multiply by the derivative of sqrt(1+x^2) which is 1/2sqrt(1+x^2)^-.5(2x) or [xsqrt(1+x^2)-.5]1/sqrt(1+x^2)]=x/x^2+1

OpenStudy (anonymous):

how about differentiate e^x/10 i know e^x is e^x

OpenStudy (mertsj):